OpenStudy (anonymous):

limit of 3x^3+ln(x)/x+3x^2 as x approaches infinity.

OpenStudy (amistre64):

maybe a salnt asymp?

OpenStudy (anonymous):

? \[\frac{3x^3+ln(x)}{x+3x^2}\]

OpenStudy (amistre64):

slant even

OpenStudy (anonymous):

saint asymptote, the patron saint of division

OpenStudy (amistre64):

otherwise i think its just inf

OpenStudy (anonymous):

\[3x^3+\ln(x)/x+2x^3\] typo'd sorry, dunno if that makes a difference.

OpenStudy (amistre64):

y = 3x might be the slant..

OpenStudy (amistre64):

it does lol

OpenStudy (amistre64):

3/2

OpenStudy (anonymous):

I'm bad at limits. Haha any help is greatly appreciated.

OpenStudy (amistre64):

the infinity limits are gona be at the horizontal asymps if they exist

OpenStudy (anonymous):

still infinity

OpenStudy (amistre64):

i think the ln(x)/x^3 goes to zero...

OpenStudy (anonymous):

oh i see it is \[\frac{3x^3+ln(x)}{x+2x^3}\] my mistake i misread it. amistre right, 3/2

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=lim {x+to+infinity}lnx%2Fx^3 im right lol

OpenStudy (anonymous):

yes you are right of course. ration of leading coefficients.

OpenStudy (anonymous):

ratio too

OpenStudy (amistre64):

ration? now you are even typing like me ;)

OpenStudy (anonymous):

How did you find that? How would you cancel out the natural log? I'm still kinda confused...

OpenStudy (amistre64):

the log tends to slow down at inifinty; so the x^3 takes over and goes to 0

OpenStudy (anonymous):

Ahh and then ratio of coefficients. Got it. Thanks!

OpenStudy (amistre64):

a the ends; log(x) is pretty much a straight line

OpenStudy (anonymous):

log grows slower than ANY polynomial, so you can ignore it safely

OpenStudy (anonymous):

Right, right, I'm in compsci. We covered that. Should have remembered. Thanks so much! This was the first question I asked, how do I close this question?

OpenStudy (amistre64):

you cant, its exists now forever in the ether of the cosmos ;)

OpenStudy (anonymous):

That's... Unfortunate.

OpenStudy (anonymous):

lord i hope not with all my wrong answers!

OpenStudy (amistre64):

if you swear alot they might delete it lol